OpenStudy (anonymous):

find the Laplace transform of g(t)= t^(3/2) ? could anyone help me with this plz?

OpenStudy (turingtest):

I did this the other day...

OpenStudy (anonymous):

i posted it yesterday and no one help me with it .

OpenStudy (turingtest):

\[L(t^{n-1/2}),n\in\mathbb N=\frac{1\cdot3\cdot5\cdot\cdot\cdot(2n-1)\sqrt\pi}{2^ns^{n+1/2}}\]in your case n=2

OpenStudy (anonymous):

SORRY 1.3.5

OpenStudy (turingtest):

you multiply all the odd numbers up through 2n-1 in your case 2n-1=4-1=3, so yours will just be\[L(t^{2-1/2})=\frac{1\cdot3\sqrt\pi}{2^2s^{2+1/2}}\]

OpenStudy (anonymous):

oh, i got it . very grateful for your help bro ..

OpenStudy (turingtest):

anytime :D